Official solution
Probablity of getting head P(H) = \[{1 \over 2}\]
Probablity of getting tail P(T) = 1 - \[{1 \over 2}\] = \[{1 \over 2}\]
Probability of observing at least one head out of n tosses
= 1 - Probability of observing no head occurs out of n tosses
= 1 - \[{\left( {{1 \over 2}} \right)^n}\]
According to the question,
1 - \[{\left( {{1 \over 2}} \right)^n}\] \[ \ge \] \[{{90} \over {100}}\]
\[ \Rightarrow \] \[{\left( {{1 \over 2}} \right)^n} \le {1 \over {10}}\]
\[ \Rightarrow \] \[{2^n} \ge 10\]
As 23 = 8
24 = 16
\[ \therefore \] Minimum value of n = 4